|Ticket Number:||20170703_2||Ticket State:||CLOSED|
|Ticket Opened:||2017-07-03 14:33||Ticket Closed:||2017-08-04 13:17|
|Ticket Description:||Scheduled Upgrade of all Optical Nodes in SWITCH backbone|
|Scheduled Starttime:||2017-07-03 10:00:00||Scheduled Endtime:||2017-07-14 14:00:00|
Problem Description:A major software upgrade of our optical equipment is planned between 03.07 and 14.07.
The upgrade is split into several time-slots during the two weeks.
We expect no impact on any services during this upgrade.
The upgrade was partially completed. Due to some of our requirements, we will have to wait for a specific version to proceed further with upgrades. This should happen no sooner than middle of September. A separate ticket will be opened for these upgrades.
Concerning this ticket:
All upgrade relevant issues were successfully tackled.
The exchange of the ROADM was completed successfully. No interruption of service was observed.
A ROADM in our pop in Cern will be replaced today around 09:00, as it was identified to be malfunctioning after the upgrade. This will interrupt several backbone connections, no services should be affected though.
All optical network elements were upgraded. During the upgrade, three minor events (ROADM reboots) with minimal impact on port flaps were observed.
On Monday July 10th we begin with one more round of upgrades, expected to end Wednesday July 12th.
For all questions about this ticket, please send mail to firstname.lastname@example.org
or call +41 44 268 15 30.